OpenStudy (anonymous):

solve the logarithmic equation: log4x + log4 (x+7)=log4(x+27)......H

OpenStudy (anonymous):

I am struggling at this point, trying to pass a pre calculus class. I need all the help I can get so that I will not be out of school. Had algebra 1 and II but lost with pre calculus... HELP!!!!

OpenStudy (anonymous):

is this the equation log4x + log4(x+7) = log4(x+27) ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

by law of logarithms log4x + log4(x+7) = log16x(x+7) so log16x(x+7 = log4(x+27) 16x(x+7) = 4(x+27) 16x^2 + 112x = 4x + 108 16x^2 + 108x - 108 = 0 simplifying 4x^2 + 27x - 27 = 0 x = 0.884 or -7.634

OpenStudy (anonymous):

I thank you very much....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hey one question.... on the question it should be log4 (x+7) = log4 (x+27).... is that the same as the problem you just worked.....

OpenStudy (anonymous):

sorry... log 4 x+log4 (x+7)= log4 (x+27)

OpenStudy (anonymous):

are you saying that these are logs to the BASE 4?

OpenStudy (anonymous):

if so then my answer is wrong

OpenStudy (anonymous):

sorry.. I went back and looked to make sure

OpenStudy (anonymous):

so it s logs to base 4?

OpenStudy (anonymous):

if so then log4 x + log4 (x+7) = log4 x(x+7) then log4 x(x+7) = log4 (x+27) and x^2 + 7x = x + 27 x^2 - x - 20 = 0 x = 4 or -5

OpenStudy (anonymous):

x = 4 since there is no real log of -5

OpenStudy (anonymous):

Ok ..sorry for the confusion... I guess I am nervous about passing this class.. It has really been stressful for me and I can't thank you enough.....

OpenStudy (anonymous):

no probs - practice these as much as you can - this will improve your confidence

OpenStudy (anonymous):

I will definitely do that...have a great day!

OpenStudy (anonymous):

ty